# On the Federated Learning Framework for Cooperative Perception

### v1

@@ Abstract
Cooperative perception is essential to enhance the efficiency and safety of future transportation systems, requiring extensive data sharing among vehicles on the road, which raises significant privacy concerns. Federated learning offers a promising solution by enabling data privacy-preserving collaborative enhancements in perception, decision-making, and planning among connected and autonomous vehicles (CAVs). However, federated learning is impeded by significant challenges arising from data heterogeneity across diverse clients, potentially diminishing model accuracy and prolonging convergence periods. This study introduces a specialized federated learning framework for CP, termed the federated dynamic weighted aggregation (FedDWA) algorithm, facilitated by dynamic adjusting loss (DALoss) function. This framework employs dynamic client weighting to direct model convergence and integrates a novel loss function that utilizes Kullback-Leibler divergence (KLD) to counteract the detrimental effects of non-independently and identically distributed (Non-IID) and unbalanced data. Utilizing the BEV transformer as the primary model, our rigorous testing on the OpenV2V dataset, augmented with FedBEVT data, demonstrates significant improvements in the average intersection over union (IoU). These results highlight the substantial potential of our federated learning framework to address data heterogeneity challenges in CP, thereby enhancing the accuracy of environmental perception models and facilitating more robust and efficient collaborative learning solutions in the transportation sector.
@@ I Introduction
Environmental perception is crucial for advanced driver assistance systems (ADAS) and autonomous driving, encompassing tasks like instance segmentation [2, 3, 1, 4], traffic scene understanding [5], 2D object detection [6], 3D object detection [7, 8, 9, 10], and multi-object tracking [11, 12, 13, 14, 15]. These technologies also play a critical role in enhancing cooperative intelligent transportation systems (C-ITS), significantly improving transportation safety and efficiency [16, 17, 18]. The rise of C-ITS has drawn considerable interest from both academia and industry, highlighting a collaborative framework that integrates various ITS components—personal, vehicle, roadside, and central systems—to surpass traditional ITS models in service quality. Connected automated vehicles (CAVs) exchange information through vehicle-to-everything (V2X) networks, including vehicle-to-vehicle (V2V), vehicle-to-infrastructure (V2I), vehicle-to-network (V2N), and infrastructure-to-network (I2N).
@@
A major challenge within C-ITS is the security and privacy of shared data. Federated learning in general offers a robust solution by enabling decentralized model training without compromising data privacy, which has already been demonstrated in many applications like wireless communications [19], Industry 4.0. [20] and robotics [21], etc. In C-ITS, this approach also allows CAVs to collaboratively participate in environment perception, traffic flow prediction, and decision-making, to improve deep neural network (DNN) models while keeping the privacy protected by avoiding exchange sensor data or latent features between CAVs [22]. Additionally, V2X communication supports seamless data exchange between vehicles, roadside units (RSUs), and cloud systems, enhancing the effectiveness of federated learning in C-ITS [23, 24, 25]. Supported by RSUs, digital twins can produce accurate labels for learning algorithms, thereby increasing the precision and performance of the systems.
@@
Despite its numerous advantages, the federated learning paradigm encounters a significant challenge in data heterogeneity when implemented across diverse clients. This issue is particularly pronounced in C-ITS perception which is also known as cooperative perception (CP) due to the varied sensor configurations among participating vehicles and infrastructure. Such variations include discrepancies in camera counts, sensor placements, and non-independently and identically distributed (Non-IID) data among traffic agents, leading to marked differences in data characteristics [26]. This diversity presents a significant challenge for traditional federated learning methods, as it may be challenging to effectively synchronize the contributions of local models towards the convergence of the global model. This paper explores strategies to harness the benefits of external data while mitigating the impacts of data heterogeneity within a federated learning framework, with the aim of significantly improving the performance of local models.
@@
To validate our approach with fair comparison, we conducted extensive experiments and follow the same BEV vision model proposed in [27], which consists of five main components: an encoder for image feature extraction; positional embedding to capture camera geometry; a cross-attention module that transitions front views to BEV; convolution layers within transformers for feature refinement; and a decoder to convert BEV representations into actionable predictions. The results demonstrate the superiority of our framework, showing improved test accuracy and reduced communication overhead across most clients, thus confirming the effectiveness of the proposed federated learning framework in addressing the complex challenges.
@@
  • To enhance training efficacy further, a dynamic adjusting loss (DALoss) function is proposed. This function is tailored to the varying data distributions among clients and the central server, fine-tuning the model’s convergence direction based on real-time data distribution insights, thus ensuring improved convergence and significantly boosting the accuracy of BEV perception outcomes.
The rest of the paper are organised as follows: Section II provides related work, discussing CP, federated learning, and the application of federated learning within the CP context. Section III details our federated learning algorithms: It begins with a formulation of the problem and describes the optimization objectives of our federated learning framework, then our proposed FedDWA algorithm and DALoss function is introduced. In Section IV, our experimental setup and datasets are explained. Our experimental results, including the comparison between our proposed algorithm with other state-of-the-art methods and corresponding ablation study are analyzed. Finally, the conclusion is drawn in Section V.
@@ II-C Federated Learning with Cooperative Perception
In CP involving CAVs and RSUs, the exchange of sensor data raises significant privacy concerns. This data, sourced from vehicles and smart city infrastructure, can reveal sensitive information about individuals, such as location and personal habits. To address this, the work [43] applied federated learning to object detection tasks in AVs, capitalizing on the method’s ability to train models without compromising data privacy. However, the broader application of federated learning in CP was initially underexplored. Recent studies such as [22] and [44] discuss the application of federated learning in CP settings. The work [45] introduces an innovative federated learning framework, H2-Fed, designed to handle Hierarchical Heterogeneity, significantly improving conventional pre-trained deep learning models for CP. Furthermore, the work [46] proposes a new client selection pipeline using V2X communications, optimizing client selection based on predicted communication latency in real CP scenarios. These developments underscore the potential for federated learning for collaborative model training in privacy-sensitive environments. Despite these advances, implementing CP through federated learning presents challenges, primarily due to heterogeneity in data distribution, computational resources, and network connectivity among devices. This variability can significantly affect learning efficiency and effectiveness. To address these challenges, the work [47] introduces a fair and efficient algorithm to manage the imbalanced distribution of data and fluctuating channel conditions. Moreover, the work [27] proposed FedBEVT, a federated transformer learning approach that focuses on BEV perception to tackle common data heterogeneity issues in CP.
In this paper, we introduce an innovative federated learning framework tailored for V2X networks in CP, specifically addressing challenges posed by heterogeneity. Our framework is distinguished by its focus on the heterogeneity caused by Non-IID data across different traffic agents. We propose a novel loss function and an advanced aggregation method designed to effectively manage the individual heterogeneity inherent in different agents, thereby enhancing the overall efficacy of model training and addressing the unique challenges of Non-IID data distribution.
@@ III-A Problem Formulation
In this paper, we establish a federated learning network consisting of $M$ clients, each equipped with a unique label $m\in\{1,2,...,M\}$. The $m$-th client contains a set of data samples denoted by $\mathbf{N}_{m}=\{(\mathbf{N}_{m,i},\mathbf{Y}_{m,i})\}_{i\in\{1,2,...,|\mathbf{N}_{m}|\}}$, where ||$\cdot$|| ⋅ | denotes the number of elements in the set, $\mathbf{N}_{m,i}$ is the set of images in the $i$-th data sample captured from a multi-view camera system, and $\mathbf{Y}_{m,i}$ is the set of corresponding BEV semantic masks. Within this framework, each client independently trains a local model using their dataset and then transmits the model updates to a central server. The server integrates these updates to improve the global model. This cycle of local training and central updating repeats across several communication rounds to refine the global model optimally.
The primary optimization objective in the federated learning framework is encapsulated by:
$L=\frac{1}{\sum_{m=1}^{M}|\mathbf{N}_{m}|}\sum_{m=1}^{M}\sum_{i=1}^{|\mathbf{N}_{m}|}L_{i},$ (1)
where $L_{i}$ denotes the loss function for the data sample, and $|\mathbf{N}_{m}|$ is the number of data samples per client. This loss function is crucial for assessing and minimizing the error throughout the distributed network of clients.
Building on the framework from [27], our model features a transformer-based BEV architecture. It comprises an image feature encoder utilizing a CNN, a BEV transformer with a multi-layer attention structure (incorporating both sparse cross-view and self-attention mechanisms), and a BEV decoder. This configuration effectively encodes images from multiple cameras and transforms them into BEV features. Processed through FedBEVT, these features are decoded to produce semantic segmentation outputs from the BEV perspective, which are essential for accurately determining the positions of surrounding vehicles on the motion plane.
Our framework introduces a novel aggregation method and a customized loss function designed to tackle two primary challenges: (i) the heterogeneity of data and distribution shifts across different clients, and (ii) the varying contributions of different clients to the convergence of the global model. These enhancements are vital to ensure robust and precise model performance in autonomous driving environments.
@@ III-B Federated Dynamic Weighted Aggregation
To address the specific challenges of applying federated learning in CP, particularly client drift due to diverse sensor types and Non-IID data distributions, we introduce the federated dynamic weighted aggregation (FedDWA) algorithm. This algorithm updates the client model described by:
@@
where $G_{m}$ represents the local model of the $m$-th client,$\eta_{m}$ is the local step-size, and $\partial_{i}(G_{m}^{-})$ denotes the gradient of the $i$-th data sample for the local model in the previous round. Each client uses a local control variable $c_{m}$ to account for variances in model updates, aligning local objectives with the global model to mitigate client-specific drifts, and $c_{m}^{-}$ denotes the local control variable in the previous round. Both $c_{m}$ and $c_{g}$ are initialized as 0.
@@ 21                  $U_{m}$ $=$ $\eta_{m}(U_{m}+c_{g}^{-}-c_{m}^{-})$
25                   send $G_{m}$, $T_{m}$ to global agent
@@ 33       $G_{g}=G_{g}^{-}+\frac{\eta_{g}}{M}\sum_{m=1}^{M}(G_{g}^{-}-G_{m})$
The KLD is calculated between the predicted data distribution of the global model, $P_{g}^{-}(\mathbf{N}_{m,i})$ from the previous round, and the distribution of $m$-th local client, $P_{m}(\mathbf{N}_{m,i})$. Where $P_{g}^{-}(\mathbf{N}_{m,i})$ is approximated by using the segmentation results of our global model and $P_{m}(\mathbf{N}_{m,i})$ is approximated by using the segmentation results of our local model.
This metric helps to properly weigh the contribution of various clients to the convergence of the server model. Each client uses gradients to update the local control variable $c_{m}$:
@@
where $\eta_{g}$, the learning rate, plays a crucial role in balancing the updates, and $G_{g}$ represents the global model. This equation captures the aggregation of differences between the global and local models, promoting the convergence of the global model.
@@ III-C Dynamic Adjusting Loss Function
The cross-entropy loss function is widely used within federated learning frameworks to train individual clients. This function effectively gauges the performance of classification models that output probabilities between 0 and 1. Nevertheless, the inherent data heterogeneity across clients, particularly with non-IID data, complicates the convergence to a global optimum.
@@
$L_{i}=\mathcal{S}\left(F(\mathbf{N}_{m,i}),\mathbf{Y}_{m,i}\right)+Q_{i},$ (10)
where $\mathcal{S}\left(F(\mathbf{N}_{m,i}),\mathbf{Y}_{m,i}\right)$ denotes the standard cross-entropy loss for the $i$-th data sample of client model $F(\mathbf{N}_{m,i})$ against the true labels $\mathbf{Y}_{m,i}$, and $Q_{i}$ is the dynamic control term defined by:
$Q_{i}=C\cdot\mathcal{D}_{\mathrm{KL}}\left(P_{g}^{-}(\mathbf{N}_{m,i}),P_{m}(\mathbf{N}_{m,i})\right)\cdot||G_{g}^{-}-G_{m}||_{2}^{2},$ (11)
where $C\cdot\mathcal{D}_{\mathrm{KL}}\left(P_{g}^{-}(\mathbf{N}_{m,i}),P_{m}(\mathbf{N}_{m,i})\right)$ scales the quadratic penalty based on the KL divergence, measuring the discrepancy between the probabilistic distributions of the global and local models from the previous communication round. $C$ is a hyper-parameter, and $||\cdot||_{2}$ denotes the $L^{2}$ norm. This divergence quantifies the differences, ensuring that local updates contribute significantly toward global model convergence.
This modification incorporates a dynamic adjustable regularization mechanism to better align each client’s model updates with the global model. This dynamic adjustment of the regularization parameter allows our approach to adapt to the evolving nature of the training process and the unique characteristics of each client. By accommodating varying client impacts and data heterogeneity, our method aims to enhance both accuracy and convergence efficiency, thereby achieving a more globally optimized model in CP applications.

### v2

@@ Abstract
Cooperative perception (CP) is essential to enhance the efficiency and safety of future transportation systems, requiring extensive data sharing among vehicles on the road, which raises significant privacy concerns. Federated learning offers a promising solution by enabling data privacy-preserving collaborative enhancements in perception, decision-making, and planning among connected and autonomous vehicles (CAVs). However, federated learning is impeded by significant challenges arising from data heterogeneity across diverse clients, potentially diminishing model accuracy and prolonging convergence periods. This study introduces a specialized federated learning framework for CP, termed the federated dynamic weighted aggregation (FedDWA) algorithm, facilitated by dynamic adjusting loss (DALoss) function. This framework employs dynamic client weighting to direct model convergence and integrates a novel loss function that utilizes Kullback-Leibler divergence (KLD) to counteract the detrimental effects of non-independently and identically distributed (Non-IID) and unbalanced data. Utilizing the BEV transformer as the primary model, our rigorous testing on FedBEVT dataset which is expanded on OpenV2V dataset, demonstrates significant improvements in the average intersection over union (IoU). These results highlight the substantial potential of our federated learning framework to address data heterogeneity challenges in CP, thereby enhancing the accuracy of environmental perception models and facilitating more robust and efficient collaborative learning solutions in the transportation sector.
@@ I Introduction
Environmental perception is crucial for advanced driver assistance systems (ADAS) and autonomous driving, encompassing tasks like instance segmentation [2, 3, 1, 4], traffic scene understanding [5], 2D object detection [6], 3D object detection [7, 8, 9, 10], and multi-object tracking [11, 12, 13, 14]. These technologies also play a critical role in enhancing cooperative intelligent transportation systems (C-ITS), significantly improving transportation safety and efficiency [15, 16, 17]. The rise of C-ITS has drawn considerable interest from both academia and industry, highlighting a collaborative framework that integrates various ITS components—personal, vehicle, roadside, and central systems—to surpass traditional ITS models in service quality. Connected automated vehicles (CAVs) exchange information through vehicle-to-everything (V2X) networks, including vehicle-to-vehicle (V2V), vehicle-to-infrastructure (V2I), vehicle-to-network (V2N), and infrastructure-to-network (I2N).
@@
A major challenge within C-ITS is the security and privacy of shared data. Federated learning in general offers a robust solution by enabling decentralized model training without compromising data privacy, which has already been demonstrated in many applications like wireless communications [18], Industry 4.0. [19] and robotics [20], etc. In C-ITS, this approach also allows CAVs to collaboratively participate in environment perception, traffic flow prediction, and decision-making, to improve deep neural network (DNN) models while keeping the privacy protected by avoiding exchange sensor data or latent features between CAVs [21, 22]. Additionally, V2X communication supports seamless data exchange between vehicles, roadside units (RSUs), and cloud systems, enhancing the effectiveness of federated learning in C-ITS [23, 24, 25]. Supported by RSUs, digital twins can produce accurate labels for learning algorithms, thereby increasing the precision and performance of the systems.
@@
Despite its numerous advantages, the federated learning paradigm encounters a significant challenge in data heterogeneity when implemented across diverse clients. This issue is particularly pronounced in C-ITS perception which is also known as cooperative perception (CP) due to the varied sensor configurations among participating vehicles and infrastructure. Such variations include discrepancies in camera counts, sensor placements, and non-independently and identically distributed (Non-IID) data among traffic agents, leading to marked differences in data characteristics [26].
In a CP situation, dealing with data heterogeneity and distribution shifts is crucial. These challenges can significantly impact the performance and reliability of the system. For example, some clients may have the same type of sensors on one hand, however, their configurations such as mounting positions, angles, or calibration may vary. On the other hand, some clients may have different types of sensors, which may lead to different data characteristics. Specifically, the dataset supplied by FedBEVT [27] incorporates data from three distinct client types: cars, buses, and trucks. Each type of vehicle has sensors installed at varying positions and angles, and the calibration among these sensors also differs, resulting in data heterogeneity. Moreover, environmental variations further influence data distribution shifts. For example, changes in lighting conditions across multiple time frames within a single scene can significantly alter, such phenomenon could be observed clearly in the dataset supplied by FedBEVT [27]. These are significant challenges for traditional federated learning methods, as it may be challenging to effectively synchronize the contributions of local models towards the convergence of the global model. This paper explores strategies to harness the benefits of external data while mitigating the impacts of data heterogeneity within a federated learning framework, with the aim of significantly improving the performance of local models.
@@
To validate our approach with fair comparison, we conducted extensive experiments and follow the same BEV semantic segmentation network model proposed in [27], which consists of five main components: an encoder for image feature extraction; positional embedding to capture camera geometry; a cross-attention module that transitions front views to BEV; convolution layers within transformers for feature refinement; and a decoder to convert BEV representations into actionable predictions for BEV semantic segmentation. The results demonstrate the superiority of our framework, showing improved test accuracy and reduced communication overhead across most clients, thus confirming the effectiveness of the proposed federated learning framework in addressing the complex challenges.
@@
  • To enhance training efficacy further, a dynamic adjusting loss (DALoss) function is proposed for CP task under federated learning framework. This function is tailored to the varying data distributions among clients and the central server, fine-tuning the model’s convergence direction based on real-time data distribution insights, thus ensuring improved convergence and significantly boosting the accuracy of BEV perception outcomes.
The rest of the paper are organised as follows: Section II provides related work, discussing CP, federated learning, and the application of federated learning within the CP context. Section III explains proposed FedDWA algorithm and DALoss function, with a formulation of the problem and describes the optimization objectives of our federated learning framework. In Section IV, our experimental setup and datasets are explained first, followed by details regarding applying FedDWA algorithm and DALoss function for BEV semantic segmentation task under CP setting. Finally, the conclusion is drawn in Section V.
@@ II-C Federated Learning with Cooperative Perception
In CP involving CAVs and RSUs, the exchange of sensor data raises significant privacy concerns. This data, sourced from vehicles and smart city infrastructure, can reveal sensitive information about individuals, such as location and personal habits. To address this, the work [43] applied federated learning to object detection tasks in AVs, capitalizing on the method’s ability to train models without compromising data privacy. However, the broader application of federated learning in CP was initially underexplored. Recent studies such as [21], [22] and [44] discuss the application of federated learning in CP settings. The work [45] introduces an innovative federated learning framework, H2-Fed, designed to handle Hierarchical Heterogeneity, significantly improving conventional pre-trained deep learning models for CP. Furthermore, the work [46] proposes a new client selection pipeline using V2X communications, optimizing client selection based on predicted communication latency in real CP scenarios. These developments underscore the potential for federated learning for collaborative model training in privacy-sensitive environments. Despite these advances, implementing CP through federated learning presents challenges, primarily due to heterogeneity in data distribution, computational resources, and network connectivity among devices. This variability can significantly affect learning efficiency and effectiveness. To address these challenges, the work [47] introduces a fair and efficient algorithm to manage the imbalanced distribution of data and fluctuating channel conditions. Moreover, the work [27] proposed FedBEVT, a federated transformer learning approach that focuses on BEV perception to tackle common data heterogeneity issues in CP.
In this paper, we introduce an innovative federated learning framework tailored for V2X networks in CP, specifically addressing challenges posed by heterogeneity. Our framework is distinguished by its focus on the heterogeneity caused by Non-IID data across different traffic agents. We propose a novel loss function and an advanced aggregation algorithm designed to effectively manage the individual heterogeneity inherent in different agents, thereby enhancing the overall efficacy of model training and addressing the unique challenges of Non-IID data distribution.
@@ III-A Problem Formulation
In this paper, we aim to establish a federated learning framework, $FL$, consisting of $M$ clients that each equipped with a unique label $m\in\{1,2,...,M\}$, for the BEV semantic segmentation task under V2X CP setting. The $m$-th client contains a set of data samples denoted by $\mathbf{N}_{m}=\{(\mathbf{N}_{m,i},\mathbf{Y}_{m,i})\}_{i\in\{1,2,...,|\mathbf{N}_{m}|\}}$, where ||$\cdot$|| ⋅ | denotes the number of elements in the set, $\mathbf{N}_{m,i}$ is the set of images in the $i$-th data sample captured from a multi-view camera system, and $\mathbf{Y}_{m,i}$ is the set of corresponding BEV semantic masks. Within this framework, each client $m$ trains a local model $G_{m}$ independently using the calculated gradient, by minimizing a primary optimization objective loss function $L_{m,i}$ for each data sample in their dataset, and then transmits the model updates to a central server. Thus this loss function is crucial for assessing and minimizing the error throughout the distributed network of clients $m$. The server integrates these updates to improve the global model. This cycle of local training and central updating repeats across several communication rounds to refine the global model optimally.
To demonstrate the effectiveness of our proposed federated learning framework, we follow FedBEVT [27] to employ the same transformer-based BEV segmentation network as network model $G_{m}$ and $G_{g}$, on each client and server respectively. Such network model comprises an image feature encoder utilizing a CNN, a BEV transformer with a multi-layer attention structure (incorporating both sparse cross-view and self-attention mechanisms), and a BEV decoder. This configuration effectively encodes images from multiple cameras and transforms them into BEV features. These features are decoded to produce semantic segmentation outputs on the BEV perspective.
Our framework introduces a novel aggregation algorithm and a customized loss function designed to tackle two primary challenges for CP setting: (i) the heterogeneity of data and distribution shifts across different clients, and (ii) the varying contributions of different clients to the convergence of the global model. These enhancements are vital to ensure robust and precise model performance in autonomous driving environments.
@@ III-B Federated Dynamic Weighted Aggregation
To address the specific challenges of applying federated learning in CP, particularly client drift due to diverse sensor types and Non-IID data distributions, we introduce the federated dynamic weighted aggregation (FedDWA) algorithm for our federated learning framework $FL$. This algorithm updates the client model described by:
@@ 21                  $U_{m}$ $=$ $\eta_{m}(U_{m}+c_{g}^{-}-c_{m}^{-})$
25                   send $G_{m}$, $T_{m}$ to global server
@@ 33       $G_{g}=G_{g}^{-}+\frac{\eta_{g}}{M}\sum_{m=1}^{M}(G_{g}^{-}-G_{m})$
where $G_{m}$ represents (the weights of) the local model of the $m$-th client, $\eta_{m}$ is the local step-size, and $\partial_{i}(G_{m}^{-})$ denotes the gradient of the $i$-th data sample for the local model in the previous communication round.
Each client uses a local control variable $c_{m}$ to account for variances in model updates, aligning local objectives with the global model to mitigate client-specific drifts, and $c_{m}^{-}$ denotes the local control variable in the previous communication round. Both $c_{m}$ and $c_{g}$ are initialized as 0.
@@
The KLD is calculated between the predicted data distribution of the global model from the previous communication round, $P_{g}^{-}(\mathbf{N}_{m,i})$, and the distribution of $m$-th local client, $P_{m}(\mathbf{N}_{m,i})$. Where $P_{g}^{-}(\mathbf{N}_{m,i})$ is approximated by using the segmentation output of global model from previous communication round $G^{-}_{g}$, and $P_{m}(\mathbf{N}_{m,i})$ is approximated by using the segmentation output of each local model $G_{m}$. To transform the segmentation output into a distribution, a log softmax function is applied to convert the segmentation output into log probabilities, e.g.:
$\begin{split}P_{m}(\mathbf{N}_{m,i})=\begin{bmatrix}log(\sigma(V_{1,1}))&\dots&log(\sigma(V_{1,k}))\\ \dots&\dots&\dots\\ log(\sigma(V_{j,1}))&\dots&log(\sigma(V_{j,k}))\end{bmatrix},\end{split}$ (6)
where $V_{j,k}$ denotes the $(j,k)$-th pixel values in the segmentation output matrix $G_{m}(\mathbf{N}_{m,i})$ and $\sigma$ represents the softmax function. The same calculation is applied for $P_{g}^{-}(\mathbf{N}_{m,i})$.
Each client uses gradients to update the local control variable $c_{m}$:
@@
where $\eta_{g}$, the learning rate, plays a crucial role in balancing the updates, and $G_{g}$ represents the global model. This equation captures the aggregation of differences between the global and local models, promoting the convergence of the global model. To provide an comprehensive understanding of FedDWA, the pseudo code is illustrated as Algorithm 1.
@@ III-C Dynamic Adjusting Loss Function
As discussed before, it is critical to define a suitable loss function $L_{m,i}$, to calculate the gradient of $i$-th data sample for each client’s local model, $\partial_{i}(G_{m})$. The cross-entropy loss function is widely used within federated learning frameworks to train individual clients. This function effectively gauges the performance of classification models that output probabilities between 0 and 1. Nevertheless, the inherent data heterogeneity across clients, particularly with non-IID data, complicates the convergence to a global optimum.
@@
$L_{m,i}=\mathcal{S}\left(G_{m}(\mathbf{N}_{m,i}),\mathbf{Y}_{m,i}\right)+Q_{m,i},$ (10)
where $\mathcal{S}\left(G_{m}(\mathbf{N}_{m,i}),\mathbf{Y}_{m,i}\right)$ denotes the standard cross-entropy loss for the $i$-th data sample of client model $G_{m}$ against the true labels $\mathbf{Y}_{m,i}$, and $Q_{m,i}$ is the dynamic control term defined by:
$Q_{m,i}=C\cdot\mathcal{D}_{\mathrm{KL}}\left(P_{g}^{-}(\mathbf{N}_{m,i}),P_{m}(\mathbf{N}_{m,i})\right)\cdot||G_{g}^{-}-G_{m}||_{2}^{2},$ (11)
where $C\cdot\mathcal{D}_{\mathrm{KL}}\left(P_{g}^{-}(\mathbf{N}_{m,i}),P_{m}(\mathbf{N}_{m,i})\right)$ scales the quadratic penalty based on the KLD, measuring the discrepancy between the probabilistic distributions of the global and local models from the previous communication round. $C$ is a regularization constant, and $||\cdot||_{2}$ denotes the $L^{2}$ norm. This divergence quantifies the differences, ensuring that local updates contribute significantly toward global model convergence.
This modification incorporates a dynamic adjustable regularization mechanism to better align each client’s model updates with the global model. Such dynamic adjustment of the regularization parameter allows our approach to adapt to the evolving nature of the training process and the unique characteristics of each client.
By integrating KLD into the loss function, the learning process is encouraged to minimize deviations between data distribution in different clients, effectively aligning the local models closer to the global model. Specifically, the KLD is employed as a weighting factor in the L2-norm calculation between the global and local models, as shown in Eq. (11). A large KLD indicates a significant discrepancy in the distributions, resulting in an increased weight. This increment in weight amplifies the loss, consequently accelerating the convergence of the local model towards the global model. By minimizing this divergence between local updates and the global model, the learning process adaptively adjusts to the unique data distributions present in each client. Therefore, our approach ensures that local updates contribute positively towards a more generalized global model, mitigating the impact of data distribution shifts and data heterogeneity.
While the design of the loss function allows for larger updates when there are significant discrepancies between the local and global models, it is crucial to manage the potential instability might be introduced by these aggressive updates. To avoid such instability, we use a regularization constant $C$, which can penalize large changes in model weights and provide a counterbalance to aggressive updates. We also use adaptive learning rates which can adjust the step size based on the training dynamics that we reduce the learning rate after several aggregation steps thus tempering the update magnitude.